Jim can row his boat upstream at four kilometers per hour, and downstream at six kilometers per hour. He rowed upstream a certain distance and then rowed back downstream to his starting point in one hour. How far did he row upstream?
Answer by Boreal(15235)

Row=4 km/h
Current=C
R-C=4 km/h
R+C=6 km/h
2R=10
R=5 km/h
C=1 km/h
t=time upstream in hr.
1-t=time downstream in hr.
4t=6(1-t)
4t=6-6t
10t=6
t=(3/5) hour or 12/5 km. upstream
1-t=(2/5) hour or 12/5 km downstream
12/5 km or 2.4 km.